Calculate the molar concentration of solute in each of the following:

  1. Exactly 7.59g of NH3 dissolved in 1.500L of solution.
  2. Exactly 0.291g of N2SO dissolved in 8.65L of solution.
  3. Exactly 85.2g of NaOH dissolved in 5.00L of solution.
  4. Exactly 31.25g of NaCl dissolved in 2.00L of solution.
  5. Exactly 6.00mg of atmospheric N2 dissolved in water in a solution volume of 2.00L.
1
Expert's answer
2021-11-04T00:32:02-0400

Molarity of "NH_3=[\frac{7.59}{17.031}]÷1.500=0.297M"


"H_2SO_4=[\frac{0.291}{
98.079 }]÷8.65=0.00034M"


"NaOH=[\frac{85.2}
{
39.997 
}]÷5.00=0.43M"


"NaCl=[\frac{31.25}{
58.44 
}]÷2.00=0.267M"